OpenStudy (anonymous):

[6.01] Select the system of equations that corresponds to the given graph. grahp.jpg 4x + 2y = 12 4x + 2y = 4 2x + 3y = 4 6x + 9y = 12 2x + 4y = 12 2x + y = 3 2x + 3y = 8 −x + y = −4

OpenStudy (anonymous):

here is the graph

jimthompson5910 (jim_thompson5910):

the two equations produce the same line (one line is right on top of the other)

jimthompson5910 (jim_thompson5910):

which system gives you something like that?

OpenStudy (anonymous):

2x + 3y = 4 6x + 9y = 12

OpenStudy (anonymous):

4x + 2y = 12 4x + 2y = 4 this one

jimthompson5910 (jim_thompson5910):

close, the left sides are identical...but the right sides are different this means you'll have 2 parallel lines

OpenStudy (anonymous):

ok? i am lost here can you explain it better to me?

OpenStudy (anonymous):

jewest19 why will it be that? i am lost?

jimthompson5910 (jim_thompson5910):

6x + 9y = 12 3(2x+3y) =3*4 2x+3y = 4 ... Divide both sides by 3. notice how this identical to the first equation in the system 2x + 3y = 4 6x + 9y = 12

jimthompson5910 (jim_thompson5910):

so that shows you that you're dealing with the same equation (which produces the same exact graph)

OpenStudy (anonymous):

Look at the graph and find any possible points you can see, (2,0) is on the graph. Plug 2 for x and 0 for y and see which equations work. The only one that works is 2x + 3y = 4 6x + 9y = 12.

OpenStudy (anonymous):

Oh i see it

OpenStudy (anonymous):

so you mean that's the answer jewest19??? or is there more???

OpenStudy (anonymous):

yes 2x + 3y = 4 6x + 9y = 12 is the answer.

OpenStudy (anonymous):

I get it now that you i am not good when it comes to graph but thank you that help me a lot so thanks.